OpenStudy (anonymous):

Determine the following Laurent Series: \[\frac{1}{z-i}+\frac{1}{z+1} \] on \[ [Z \in C, 0<\left| Z +1 \right|<\sqrt{2}]\]

OpenStudy (anonymous):

Use the geometric series..

OpenStudy (anonymous):

\[\frac{1}{z+1-(1+i)}\]

OpenStudy (anonymous):

?? i was thinking something like this

OpenStudy (anonymous):

\[-\frac{1}{1-i} \frac{1}{1-\frac{z+1}{1+i}}\]

OpenStudy (anonymous):

It depends on the point around which you want to develop the function.

OpenStudy (anonymous):

sorry, I mean the expansion point, English is not my native language, I hope you know what I mean